# Topological properties of curved spacetime Su–Schrieffer–Heeger model

### v2

@@ Abstract
The Su-Schrieffer-Heeger (SSH) model, a prime example of a one-dimensional topologically nontrivial insulator, has been extensively studied in flat space-time. In recent times, many studies have been conducted to understand the properties of the low-dimensional quantum matter in curved spacetime, which can mimic the gravitational event horizon and black hole physics. However, the impact of curved spacetime on the topological properties of such systems remains unexplored. Here, we investigate the curved spacetime (CST) version of the SSH model by introducing a position-dependent hopping parameter. We show, using different topological markers, that the CST-SSH model can undergo a topological phase transition. We find that the topologically non-trivial phase can host zero-energy edge modes, but those edge modes are asymmetric, unlike the usual SSH model. Moreover, we find that at the topological transition point, a critical slowdown takes place for zero-energy wave packets near the boundary, indicating the presence of a horizon, and interestingly, if one moves even a slight distance away from the transition point, wave packets start bouncing back and reverse direction before reaching the horizon. An extended version of the CST-SSH model, which supports distinct topological phases characterized by higher winding numbers, has also been studied. We show that a similar critical slowdown of wave packets near the boundary occurs at all topological transition points between any two distinct phases. A semiclassical description of the wave packet trajectories also supports these results.
@@ I Introduction
In this work, our main goal is to connect these two extremely different branches of physics, 1) curved spacetime and black hole, and 2) topological systems in condensed matter. We propose a CST lattice model inspired by the two-band SSH model. First, we investigate whether we can see the event horizon physics or not. Next, we investigate whether the CST version of the SSH model has a topologically non-trivial phase or not. Remarkably, we answer both questions affirmatively in this manuscript. We show using different topological markers that the CST-SSH model also shows a topological phase transition. Most interestingly, exactly at the transition point, a critical slowdown takes place for zero-energy wave packets, which indicates the presence of a horizon. The same phenomenon has also been observed in the extended version of the CST-SSH model at all topological transition points.
This paper is organized as follows: We define our CST-SSH model in Section II. Section III shows the wave packet dynamics and semi-classical analysis of the dynamics. We investigate the gap of the spectrum and also the zero-energy states in Sec. IV. We dedicate the Sec. V for different topological markers to check the topological nature of the model, and Sec. VI shows the effect of quench across different topological phases in CST-SSH. Section . VII is dedicated for the extended version of the CST-SSH model. Finally, we summarize our findings in Sec. VIII.
@@ II Curved spacetime SSH models
where $c^{\dagger}_{n}$ and $c_{n}$ are fermionic-creation and annihilation operators, and $t_{n}$ is a position-dependent hopping parameter. In the thermodynamic limit, if $t_{n}=t$ (the nn hopping amplitude is constant for all sites), then the Hamiltonian can be diagonalized very easily in the momentum space $k$. The Hamiltonian will read as, $\mathcal{H}=\sum_{k}\epsilon(k)c^{{\dagger}}_{k}c_{k}$, where $c^{\dagger}_{k}$ and $c_{k}$ are creation and annihilation operators in momentum space, and $\epsilon(k)=-2t\cos k$. In case of position-dependent hopping $t_{n}$, with $N\rightarrow\infty$, we can approximate the local band structure for the bulk as $\epsilon(n,k)\sim-2t_{n}\cos k$, and the local Fermi velocity as $v_{F}(n)\sim\pm 2t_{n}$. Thus, the local velocity varies with spatial coordinates, as the hopping here is position-dependent. In the continuum limit, these kinds of lattice models are equivalent to a massless Dirac field with the two-component spinor $\hat{\Psi}=(\hat{\psi}_{+},\hat{\psi}_{-})^{T}$ that obeys the following time-evolution equation [46],
@@
At $v(x)=0$, i.e., $t(x)=0$, this Rindler metric possesses an event horizon, as at that point, the local speed of light goes to zero. If we let $t(x)\propto x^{\sigma}$, where $\sigma$ is indicating the warping of spacetime [22], at $x=0$, the lattice model will possess event horizon for $\sigma\geq 1$: an gaussian wave-packet propagating in this lattice will suffer eternal slowdown and never reaches the origin, thus effectively forming an event horizon at the origin in the lattice model.
In a similar spirit, in this work, we study the curved-space-time version of the SSH model, which is described by the following Hamiltonian,
$H=\sum_{n=1}^{N}t_{1}(n)c^{\dagger}_{n,A}c_{n,B}+\sum_{n=1}^{N-1}t_{2}(n)c^{\dagger}_{n,B}c_{n+1,A}+h.c.$ (4)
where $c^{\dagger}_{n,A}$, $c_{n,A}$ ($c^{\dagger}_{n,B}$, $c_{n,B}$) are fermionic creation and annihilation operators at the A sub-lattice (B sub-lattice) of the nth unit cell and $t_{1}(n)=t_{1}\big{(}\frac{2n-1}{2N-1}\big{)}^{\sigma}$ and $t_{2}(n)=t_{2}\big{(}\frac{2n}{2N-1}\big{)}^{\sigma}$.
In the $\sigma=0$ limit, the model maps exactly to the usual SSH model, which has a topologically non-trivial (trivial) phase when $t_{1}<t_{2}$ ($t_{1}>t_{2}$), and is followed by a topological transition when $t_{1}=t_{2}$ [40]. We identify $t_{1}$ and $t_{2}$ as the intra- and inter-cell hopping strengths, respectively.
@@
We also consider an extended version of the SSH model, where the hopping is not just restricted to between nearest neighbors, but higher-order hopping has also been allowed. In the context of the standard SSH model, a similar model was studied in Ref. [47, 48], which is known as the extended SSH model. Here we study the CST version of such models, where hopping is position-dependent, and the model reads as,
$H_{ext}$ $=\sum_{n=1}^{N}\Bigl{(}t_{1}(n)\,c^{\dagger}_{n,A}c_{n,B}\;+\;t_{2}(n)\,c^{\dagger}_{n,B}c_{n+1,A}$ (5)
$\quad+t_{3}(n)\,c^{\dagger}_{n,A}c_{n+1,B}\;+\;t_{4}(n)\,c^{\dagger}_{n,B}c_{n+2,A}+\mathrm{H.c.}\Bigr{)},$
@@
where, $t_{1}(n)=t_{1}\big{(}\frac{2n-1}{2N-1}\big{)}^{\sigma}$, $t_{2}(n)=t_{2}\big{(}\frac{2n}{2N-1}\big{)}^{\sigma}$ are the nearest neighbor intracellular and intercellular position-dependent hoppings, and, $t_{3}(n)=t_{3}\big{(}\frac{2n-1}{2N-1}\big{)}^{\sigma}$, $t_{4}(n)=t_{4}\big{(}\frac{2n}{2N-1}\big{)}^{\sigma}$. For $\sigma=0$, the Hamiltonian will have translational invariance. Hence, it can be fourier transformed to momentum $k$ space, where the winding number transition can happen when the components of the bloch vector becomes zero, i.e., $d_{x}(k)=t_{1}+(t_{2}+t_{3})\cos k+t_{4}\cos 2k$ and, $d_{y}(k)=(t_{2}-t_{3})\sin k+t_{4}\sin 2k$ both becomes 0, which gives us conditions: $(t_{1}+t_{2}+t_{3}+t_{4}=0)$, $(t_{1}-(t_{2}+t_{3})+t_{4}=0)$ , and, $(t_{1}=t_{4}$ and $t_{2}=t_{3})$ at $k=0,~\pi$~ and $\pi/2$, respectively, where we can observe winding number transitions.
Figure 1: Time evolution of Gaussian wave packet in the CST-SSH Hamiltonian of system size $2N=1000$ (500 unit cells): (a) $t_{1}=t_{2}=1,\sigma=1,\omega=25,x_{0}=750,p_{0}=-\frac{\pi}{2}$, the wave packet experiences an eternal slowdown as it approaches the origin; (b) $t_{1}=1,t_{2}=1.9,\sigma=1,\omega=25,x_{0}=750,p_{0}=-\frac{\pi}{2}$, the wave-packet turns back before the origin; (c) peak position of the Gaussian wave packet vs. time for different $p_{0}$, at $\sigma=1$, having $(t_{1},t_{2})=(1,1.9)$; (d) peak position vs. time for different $t_{1}/t_{2}$ ratios at $p_{0}=-\frac{\pi}{2}$. Dotted lines are numerical; solid lines are semiclassical predictions.
0.2
@@ III Wave-packet vs semiclassical dynamics
which describes a normalized gaussian wave packet at time $\tau=0$ centered at position $x_{0}$ of the lattice, and having initial momentum of $p_{0}$, $\omega$ is the width of the wave packet. We find that if $t_{1}\neq t_{2}$, the wave-packet first propagates towards an edge and then returns from some point $x_{min}$, which we identify as a ‘turning point’. This $x_{min}$ depends on the details, e.g., initial momentum $p_{0}$, the Hamiltonian parameters $\sigma$, $t_{1}$, and $t_{2}$ (see Fig. 1b). On the other hand, for $t_{1}=t_{2}$ and $p_{0}=-\pi/2$, it eternally slows down while evolving as it approaches the edge $x=0$, and never returns as shown in Fig. 1a. We can identify it as the asymptotic localization of wave packets at the origin, which mimics the key feature expected for wave packet dynamics in the presence of a horizon.
We can dive deeper into the time evolution of the wave packet by obtaining the semiclassical trajectories of the Hamiltonian. For a position-dependent SSH model, the two recursive energy eigenvalue equations can be written as,
$t_{2}(n-1)\psi_{B,n-1}+t_{1}(n)\psi_{B,n}=\epsilon\psi_{A,n}$ (7)
$t_{1}(n)\psi_{A,n}+t_{2}(n)\psi_{A,n+1}=\epsilon\psi_{B,n}$ (8)
@@
$\Big{(}t_{2}(\hat{x}-2\delta x)e^{-2i\delta x\hat{p}}+t_{1}(\hat{x})\Big{)}\tilde{\psi}_{B}(x,\tau)=i\partial_{\tau}\tilde{\psi}_{A}(x,\tau)$ (9)
@@
$\Big{(}t_{1}(\hat{x})+t_{2}(\hat{x})e^{2i\delta x\hat{p}}\Big{)}\tilde{\psi}_{A}(x,\tau)=i\partial_{\tau}\tilde{\psi}_{B}(x,\tau),$ (10)
@@
where, $\tilde{\psi}_{A/B}(x,\tau)=\tilde{\psi}_{A/B}(x)e^{-i\epsilon\tau}$. Combining the two equations, we get the continuum Hamiltonian as a $2\times 2$ matrix,
$\tilde{H}=\begin{pmatrix}0&t_{1}(\hat{x})+e^{-2i\,\delta x\,\hat{p}}\,t_{2}(\hat{x})\\[2.84526pt] t_{1}(\hat{x})+t_{2}(\hat{x})\,e^{2i\,\delta x\,\hat{p}}&0\end{pmatrix}$,
which acts on the spinor $\Psi(x,\tau)=\bigl{(}\begin{smallmatrix}\tilde{\psi}_{A}(x,\tau)\\ \tilde{\psi}_{B}(x,\tau)\end{smallmatrix}\bigr{)}$. Now, neglecting the commutation relation between $\hat{x}$ and $\hat{p}$, and then, rescaling the momentum and time as $p\rightarrow p/\delta x$ and $\tau\rightarrow\tau/\delta x$, we could express the energy as follows, and with it, we can obtain the equation of motion. The semiclassical equations of motion (in one dimension) read:
$\dot{x}\;=\;\frac{\partial E_{\pm}}{\partial p},\qquad\dot{p}\;=\;-\,\frac{\partial E_{\pm}}{\partial x},$ (11)
where,
$E_{\pm}(x,p)\;=\;\pm\,\sqrt{\,t_{1}^{2}(x)\;+\;t_{2}^{2}(x)\;+\;2\,t_{1}(x)\,t_{2}(x)\,\cos\!\bigl{(}\,2p\bigr{)}}$ (12)
is the energy expression, and correspondingly, one can obtain equations of motion as,
$\dot{x}\;=\;\frac{\partial E_{\pm}}{\partial p}\;=\;\mp\,\frac{\,2t_{1}(x)\,t_{2}(x)\sin\bigl(2p\bigr{missing})\,}{\sqrt{\,t_{1}^{2}(x)+t_{2}^{2}(x)+2\,t_{1}(x)\,t_{2}(x)\,\cos(2p)\,}}$ (13)
@@
Solving these two coupled differential equations, one can easily obtain the semiclassical trajectories. Here, the $\mp$ signs just indicate the direction of the propagation of the trajectories. In figure 1c and 1d, the semi-classical trajectories are compared with the peak position of exact dynamics results. The results show extraordinary agreement. Moreover, the semiclassical result can also mimic the event horizon physics for $t_{1}=t_{2}$ and $p_{0}=-\pi/2$.
Semiclassical trajectories can also allow us to predict the turning point for $p_{0}\neq-\pi/2$. One can analytically obtain the turning point by taking the conserved nature of the average energy of wave packet from the semiclassical calculation. From equation 12, on obtains,
$\begin{split}&E^{2}(x,p)=E^{2}_{0}(x_{0},p_{0})=t_{1}^{2}(x)\;+\;t_{2}^{2}(x)\;+\;2\,t_{1}(x)\,t_{2}(x)\,\cos\!\bigl{(}\,2p\bigr{)}\\ &\implies\cos(2p)=\frac{E_{0}^{2}-(t_{1}^{2}(x)\;+\;t_{2}^{2}(x))}{2\,t_{1}(x)\,t_{2}(x)}.\end{split}$ (15)
Now, at the turning point $(x_{\min},t_{\min})$, $\dot{x}=0$, which implies $p=0$, substituting which in Eq. 15, one gets,
$E_{0}=t_{1}(x_{\min})+t_{2}(x_{\min})$ (16)
with the position dependent hoppings $t_{1}(x_{\min})\sim t_{1}x_{\min}^{\sigma}$ and $t_{2}x_{\min})\sim t_{2}x_{\min}^{\sigma}$. The Eq.  16 can be simplified to obtain the $x_{\min}$ expression as,
@@
Figure. 2 shows how the turning point varies with different $p_{0}$ values at different $\sigma$. The numerical and analytical turning points are extremely well lined up, solidifying the validity of semiclassical calculations. Moreover, the equation suggests that for $x_{\min}$ to be zero, the two hopping $t_{1}$ and $t_{2}$ must be equal, leading to a more simplified expression $x_{min}=x_{0}(\cos p)^{\frac{1}{\sigma}}$, when $p=\pm\pi/2$, only then $x_{\min}$ becomes 0. Hence, in the case of $t_{1}\neq t_{2}$, the wave packet has no initial momentum $p_{0}$ value for which the turning point becomes 0. It also validates our previous numerical finding, i.e., only when $t_{1}=t_{2}$, the model can show the synthetic horizon effect, and that too for $p_{0}=-\pi/2$, which also has been observed for the tight binding lattice model [25, 46, 22].
Figure 2: For the CST-SSH model having system size $2N=1000$ (500 unit cells), $t_{1}=1,t_{2}=1.5,$ plotting of turning point by numerical calculation ( black dotted lines) and from semiclassical formula 17 (colored dashed lines) at different initial momenta $p_{0}$ for different values of $\sigma$
Figure 3: Closing of the energy band gap and variation of zero energy eigen states variation with $\sigma$: for a CST-SSH model having system size $2N=1000$ (500 unit cells) with $t_{1}=0.5$ and $t_{2}=1$. Top row (a)–(c): Eigenvalue spectra for $\sigma=0,0.5$, and 11 respectively. Bottom row (d)–(f): Corresponding zero-energy eigenstates profiles.

### v3

# Topological properties of curved spacetime extended Su–Schrieffer–Heeger model

@@ Abstract
The Su-Schrieffer-Heeger (SSH) model, a prime example of a one-dimensional topologically nontrivial insulator, has been extensively studied in flat space-time. In recent times, many studies have been conducted to understand the properties of the low-dimensional quantum matter in curved spacetime, which can mimic the gravitational event horizon and black hole physics. However, the impact of curved spacetime on the topological properties of such systems remains unexplored. Here, we investigate the curved spacetime (CST) version of the extended SSH model, by introducing a position-dependent hopping parameter. The extended SSH model already exhibits topological phases and the associated phase transitions. Different topological markers suggest that for the same choice of parameters, the CST version of the model retains the imprint of the same topological phases and transitions. Furthermore, the topologically non-trivial phase of the CST model hosts zero-energy edge modes, which are spatially asymmetric in contrast to those of the conventional SSH model. We find that at the topological transition points between phases with different winding numbers, a critical slowdown takes place for zero-energy wave packets near the boundary, indicating the presence of a horizon, and interestingly, if one moves even a slight distance away from the topological transition points, wave packets start bouncing back and reverse direction before reaching the horizon. Moreover, we have also quantified the time scale of the critical slowdown of the wavepacket across different winding-number transition phases. A semiclassical description of the wave packet trajectories also supports these results.
@@ I Introduction
In this work, our main goal is to connect these two extremely different branches of physics, 1) curved spacetime and black hole, and 2) topological systems in condensed matter. We propose a CST lattice model inspired by the two-band extended SSH models. First, we investigate whether we can see the event horizon physics or not. Next, we investigate whether the CST version of the extended SSH models has a topologically non-trivial phase or not. Remarkably, we answer both questions affirmatively in this manuscript. We show using different topological markers that the CST-SSH model and its extended version also show a topological phase transition. Most interestingly, exactly at the transition point, a critical slowdown takes place for zero-energy wave packets, which indicates the presence of a horizon.
This paper is organized as follows: We define our CST-SSH models in Section II. Section III shows the wave packet dynamics and semi-classical analysis of the dynamics.We investigate the gap of the spectrum and also the zero-energy states in Sec. IV. We dedicate the Sec. V for different topological markers to check the topological nature of the models. Section. VI shows the effect of quench across different topological phases in CST-SSH. We then quantify the slowdown parameter in Sec. VII for the extended CST-SSH. Finally, we summarize our findings in Sec. VIII.
@@ II Curved spacetime SSH models
where $c^{\dagger}_{n}$ and $c_{n}$ are fermionic-creation and annihilation operators, and $t_{n}$ is a position-dependent hopping parameter. If $t_{n}=t$ (the nn hopping amplitude is constant for all sites), then the Hamiltonian can be diagonalized very easily in the momentum space $k$. The Hamiltonian will read as, $\mathcal{H}=\sum_{k}\epsilon(k)c^{{\dagger}}_{k}c_{k}$, where $c^{\dagger}_{k}$ and $c_{k}$ are creation and annihilation operators in momentum space, and $\epsilon(k)=-2t\cos k$. In case of position-dependent hopping $t_{n}(~\approx t\big(\frac{n}{N}\big)^{\sigma})$, and in the limit $N\rightarrow\infty$, the difference between the two adjacent hopping amplitudes, $t_{n+1}-t_{n}=t\Big[(x+\frac{1}{N})^{\sigma}-(x)^{\sigma}\Big]$, where, $x=\frac{n}{N}$. Now, the 1st term, in the limit $Nx$ large, read as $\Big(x+\frac{1}{N}\Big)^{\sigma}=x^{\sigma}\Big(1+\frac{\sigma}{Nx}+\mathcal{O}((\frac{1}{Nx})^{2})\Big)\text{. Therefore, }(t_{n+1}-t_{n})\approx t\Big(\frac{x^{\sigma-1}\sigma}{N}\Big)$ . Thus, in the thermodynamic limit, $N\xrightarrow{}\infty$ and for the bulk, i.e., $n\gg 0$, the hopping variation in neighboring sites is effectively negligible, i.e. for a short segment of the lattice, the hopping is effectively constant. Thus, we can approximate the local band structure for the bulk as $\epsilon(n,k)\sim-2t_{n}\cos k$, and the local Fermi velocity as $v_{F}(n)\sim\pm 2t_{n}$. Thus, the local velocity varies with spatial coordinates, as the hopping here is position-dependent. In the continuum limit, these kinds of lattice models are equivalent to a massless Dirac field with the two-component spinor $\hat{\Psi}=(\hat{\psi}_{+},\hat{\psi}_{-})^{T}$ that obeys the following time-evolution equation [38],
@@
At $v(x)=0$, i.e., $t(x)=0$, this Rindler metric possesses an event horizon, as at that point, the local speed of light goes to zero. If we let $t(x)\propto x^{\sigma}$, where $\sigma$ is indicating the warping of spacetime [31], equivalently, in the lattice model  (1) $t_{n}=(\frac{n}{N-1})^{\sigma}$, at $n/N\to 0$ limit, the lattice model will possess event horizon for $\sigma\geq 1$: an gaussian wave-packet propagating in this lattice will suffer eternal slowdown and never reaches the origin, thus effectively forming an event horizon at the origin in the lattice model.
@@
In a similar spirit, in this work, we investigate whether we observe the horizon physics in a two-band topological system with next nearest neighbor hopping, and what would happen to the topological characteristics after introducing position dependent hopping. Hence, we study an extended version of the SSH model, where the hopping is not just restricted to between nearest neighbors, but higher-order hopping has also been allowed. In the context of the standard SSH model, a similar model was studied in Ref. [61, 22], known as the extended SSH model. Here we study the CST version of such models, where hopping is position-dependent, and the model reads as,
$H$ $=\sum_{n=1}^{N}t_{1}(n)\,c^{\dagger}_{n,A}c_{n,B}\;+\;\sum_{n=1}^{N-1}t_{2}(n)\,c^{\dagger}_{n,B}c_{n+1,A}$ (4)
$\quad+t_{3}(n)\,c^{\dagger}_{n,A}c_{n+1,B}\;+\;\sum_{n=1}^{N-2}t_{4}(n)\,c^{\dagger}_{n,B}c_{n+2,A}+h.c.,$
@@
where $c^{\dagger}_{n,A}$, $c_{n,A}$ ($c^{\dagger}_{n,B}$, $c_{n,B}$) are fermionic creation and annihilation operators at the A sub-lattice (B sub-lattice) of the nth unit cell, and, $t_{1}(n)=t_{1}\big(\frac{2n-1}{2N-1}\big)^{\sigma}$, $t_{2}(n)=t_{2}\big(\frac{2n}{2N-1}\big)^{\sigma}$ are the nearest neighbor intracellular and intercellular position-dependent hoppings, and, $t_{3}(n)=t_{3}\big(\frac{2n-1}{2N-1}\big)^{\sigma}$, $t_{4}(n)=t_{4}\big(\frac{2n}{2N-1}\big)^{\sigma}$. For $\sigma=0$, the Hamiltonian will have translational invariance. Hence, it can be fourier transformed to momentum $k$ space, where the winding number transition can happen when the components of the bloch vector becomes zero, i.e., $d_{x}(k)=t_{1}+(t_{2}+t_{3})\cos k+t_{4}\cos 2k$ and, $d_{y}(k)=(t_{2}-t_{3})\sin k+t_{4}\sin 2k$ both becomes 0, which gives us conditions:
$t_{1}+t_{2}+t_{3}+t_{4}$ $=0,$ (5)
$t_{1}-(t_{2}+t_{3})+t_{4}$ $=0,$
$t_{4}(t_{1}-t_{4})+t_{3}(t_{3}-t_{2})$ $=0.$
at gap closing momentum $k_{c}=0,~\pi$~ and for $k_{c}$ satisfying $\cos k_{c}=\frac{t_{3}-t_{2}}{2t_{4}};$ ($t_{4}\neq 0,|\frac{t_{3}-t_{2}}{2t_{4}}|\leq 1)$, respectively, where we can observe winding number transitions. To gain insight into the topological properties , we also consider a limiting case of the Hamiltonian 4, with $t_{3}=t_{4}=0$, which can be considered as the curved-space-time version of the standard SSH model.
@@ III Wave-packet vs semiclassical dynamics
which describes a normalized gaussian wave packet at time $\tau=0$ centered at position $x_{0}$ of the lattice, and having initial momentum of $p_{0}$, $\omega$ is the width of the wave packet. We observe that, for the parameter values satisfying any of the conditions at Eq. 5, the wavepacket having the initial momentum of $p_{0}=k_{c}/2$ suffers a critical slowdown down while approaching towards the origin, and , never returns (Figs.1a, 1b), for $\sigma\geq 1$. On the other hand, for any other combinations of $(t_{1},t_{2},t_{3},t_{4})$, it comes back before reaching the origin 1c, which we identify as a ‘turning point’. This $x_{\min}$ depends on the details, e.g., initial momentum $p_{0}$, the Hamiltonian parameters $\sigma$, $t_{1},t_{2},t_{3}$, and $t_{4}$. In case of the CST version of the standard SSH, we find that if $t_{1}\neq t_{2}$, the wave-packet first propagates towards an edge and then returns from some point $x_{\min}$, and, for $t_{1}=t_{2}$ and $p_{0}=-\pi/2$, it eternally slows down while evolving as it approaches the edge $x=0$, and never returns. We can identify it as the asymptotic localization of wave packets at the origin, which mimics the key feature expected for wave packet dynamics in the presence of an event horizon.
[…]
III.1 Semiclassical dynamics for the Extended CST-SSH model
We can dive deeper into the time evolution of the wave packet by obtaining the semiclassical trajectories of the Hamiltonian of the extended version of the CST-SSH model. For a position-dependent extended SSH model, the two recursive energy eigenvalue equations can be written as,
$=$ $\epsilon\psi_{A,n}$
@@ III.1 Semiclassical dynamics for the Extended CST-SSH model
$\Bigl($ $t_{4}(\hat{x}-4\delta x)e^{-4i\delta x\hat{p}}+t_{3}(\hat{x})e^{2i\delta x\hat{p}}$ (9)
$+t_{2}(\hat{x}-2\delta x)e^{-2i\delta x\hat{p}}+t_{1}(\hat{x})\Bigr)\tilde{\psi}_{B}(x,\tau)=i\,\partial_{\tau}\tilde{\psi}_{A}(x,\tau)$
@@
$\Bigl(t_{4}(\hat{x})e^{4i\delta x\hat{p}}$ $+t_{3}(\hat{x}-2\delta x)e^{-2i\delta x\hat{p}}$ (10)
$+t_{2}(\hat{x})e^{2i\delta x\hat{p}}+t_{1}(\hat{x})\Bigr)\tilde{\psi}_{A}(x,\tau)=i\,\partial_{\tau}\tilde{\psi}_{B}(x,\tau)$
@@
where, $\tilde{\psi}_{A/B}(x,\tau)=\tilde{\psi}_{A/B}(x)e^{-i\epsilon\tau}$. Combining the two equations, we get the continuum Hamiltonian as a $2\times 2$ matrix
$\tilde{H}=\begin{pmatrix}0&T(\hat{x},\hat{p})\\[2.84526pt] T^{*}(\hat{x},\hat{p})&0\end{pmatrix}$,
[…]
Here, The energy expression,
$E_{\pm}(x,p)=\pm|T(x,p)|=\pm\sqrt{T_{R}^{2}(x,p)+T_{I}^{2}(x,p)},$
where, $T_{R}=t_{1}(x)+(t_{2}(x)+t_{3}(x))\text{cos}(2p)+t_{4}(x)\text{cos}(4p)$ and $T_{I}(x,p)=(t_{3}(x)-t_{2}(x))\text{sin}(2p)-t_{4}(x)\text{sin}(4p)$ are the real and imaginary part of $T(x,p)$. Now, after substituting the values of $T_{R}$ and $T_{I}$, one can obtain the equations of motion as:
where, $\partial_{p}$ and $\partial_{x}$ are the partial derivatives with respect to $p$ and $x$, respectively. Here, the $\mp$ signs just indicate the direction of the propagation of the trajectories. Fig. 1d shows the comparison between the numerical time evolution ( dotted black lines) and the semiclassical equations of motion ( colored lines ) to obtain the peak positions of the wave packets at various times for different hopping parameters. The semiclassical calculations show excellent agreement with the numerical time evolution of the Gaussian wavepackets. Semiclassical trajectories can also allow us to predict the turning point. One can analytically obtain the turning point by taking the conserved nature of the average energy of wave packet from the semiclassical calculation. One can obtain the analytical value of the turning point by equating the energy expression at $(x_{0},p_{0})$ and at $(x_{\min},p)$, where choosing an appropriate $p$ makes $\dot{x}=0$. For $p=0$, the turning position reads as:
$x_{\min}=x_{0}\Biggl[\frac{(t_{1}+(t_{2}+t_{3})\cos 2p_{0}+t_{4}\cos 4p_{0})^{2}+((t_{3}-t_{2})\sin 2p_{0}-t_{4}\sin 4p_{0})^{2}}{(t_{1}+t_{2}+t_{3}+t_{4})^{2}}\Biggr]^{\frac{1}{2\sigma}}$ (14)
In Eq. (14), the expression in the numerator is exactly the squared band energy of the corresponding homogeneous extended SSH chain evaluated at momentum $p_{0}$, while the denominator is the squared band energy at $k=0$. Equation (14) has been derived by equating the energy at $(x_{0},p_{0})$ and at $(x_{\min},p)$, and then choosing $p=0$ so that the semiclassical velocity $\dot{x}=\partial_{p}E(x,p)$ vanishes at the turning point. We have also assumed $(t_{1}+t_{2}+t_{3}+t_{4})\neq 0$, so that the denominator remains finite and the turning point is determined entirely by the value of the numerator. For fixed hopping amplitudes and fixed $x_{0}$, the turning position $x_{\min}$ is therefore controlled by how large the homogeneous band energy at the initial momentum $p_{0}$ is: when this energy decreases, the numerator of Eq. (14) becomes small and the packet penetrates deeper; when the numerator vanishes, $x_{\min}$ becomes zero.
The condition that the numerator of Eq. (14) is zero is precisely the bulk gap closing condition of the homogeneous extended SSH model evaluated at the initial momentum $p_{0}$. This can be seen explicitly in, for example, at $|p_{0}|=\pi/2$, we have the numerator becomes $\big(t_{1}-(t_{2}+t_{3})+t_{4}\big)^{2}$ , and $x_{\min}=0$ occurs when $t_{1}-(t_{2}+t_{3})+t_{4}=0$, which is the band closing at $k=\pi$. Finally, for a generic critical momentum $p_{0}=p_{c}$, the condition $x_{\min}=0$ requires the numerator of Eq. (30) to vanish. Since this numerator is a sum of two squares, both terms must vanish simultaneously. This yields $t_{4}\,(t_{1}-t_{4})+t_{3}\,(t_{3}-t_{2})=0$, with $\cos(2p_{c})=\frac{t_{3}-t_{2}}{2t_{4}}$, which is precisely the finite-momentum gap-closing condition of the extended SSH chain, with $k_{c}=2p_{c}$.
More generally, Eq. (14) is just one branch of the turning point obtained by choosing $p=0$ at the turning position. The semiclassical condition $\dot{x}=\partial_{p}E(x,p)=0$ can, however, be satisfied at other values of $p$ as well. For example, at $p=\pi/2$ the group velocity of the homogeneous extended SSH band also vanishes, and repeating the same steps gives an analogous expression for $x_{\min}$ in which the denominator is the squared band energy at $k=\pi$, namely $(t_{1}-t_{2}-t_{3}+t_{4})^{2}$, while the numerator remains the same function of $p_{0}$ as in Eq. (14), which one might use in case of $(t_{1}+t_{2}+t_{3}+t_{4})=0$. at $p_{0}=0$, in this expression, the numerator reduces to $(t_{1}+t_{2}+t_{3}+t_{4})^{2}$, and we obtain $x_{\min}=0$ when $(t_{1}+t_{2}+t_{3}+t_{4})=0$, which is the band closing at $k=0$. Thus, all the bulk gap closing conditions at Eq.5 of the extended SSH model can be recovered consistently from the turning point expression by demanding that the numerator of Eq. (14) vanishes at suitable initial momenta $p_{0}$. In the present work, we use Eq. (14) as a convenient representative branch to illustrate how the extended CST–SSH dynamics encodes the three bulk gap closing constraints through the dependence of $x_{\min}$ on the hopping amplitudes and the initial momentum $p_{0}$.
Thus, we find that the gap-closing conditions for homogeneous extended SSH models, which determine the topological transitions between different winding numbers, also correspond to the conditions under which $x_{\min}$ becomes zero in the extended CST-SSH model. This indicates that horizon physics in the extended CST-SSH model emerges only at those parameter values where topological phase transitions occur in the homogeneous extended SSH model.
Figure 2: Plotting of turning point by numerical calculation ( black dotted lines) and from semiclassical formula 17 (colored dashed lines) at different initial momenta $p_{0}$ for different values of $\sigma$, at $(t_{1},t_{2},t_{3},t_{4})=(1,1.5,0,0)$, for system size $2N=1000$ (500 unit cells),
Figure 3: Closing of the energy band gap and variation of zero energy eigen states variation with $\sigma$: for a system size $2N=1000$ (500 unit cells) with $(t_{1},t_{2},t_{3},t_{4})=(0.5,1,0,0)$ . Top row (a)–(c): Eigenvalue spectra for $\sigma=0,0.5$, and 11 respectively. Bottom row (d)–(f): Corresponding zero-energy eigenstates profiles.
In a similar spirit to the previous calculations, we can analyze the semiclassical dynamics for the CST version of the standard SSH Hamiltonian, the limiting case of Eq. 4 with $t_{3}=t_{4}=0$. After obtaining the recursive energy eigenvalue equations and the continuum Hamiltonian, we can obtain the semiclassical equation of motion as:
$\dot{x}\;=\;\frac{\partial E_{\pm}}{\partial p}\;=\;\mp\,\frac{\,2t_{1}(x)\,t_{2}(x)\sin\bigl(2p\bigr.)\,}{\sqrt{\,t_{1}^{2}(x)+t_{2}^{2}(x)+2\,t_{1}(x)\,t_{2}(x)\,\cos(2p)\,}}$ (15)
@@
In Figs.  1e and 1f, the semi-classical trajectories are compared with the peak position of exact dynamics results. Moreover, the semiclassical result can also mimic the event horizon physics for $t_{1}=t_{2}$ and $p_{0}=-\pi/2$. With $t_{1}=t_{2}$, the Hamiltonian matrix essentially becomes identical with the CST-TB Hamiltonian [40], in other words, it can also be referred to as the topological transition point of the usual SSH model. Semiclassical equations can similarly allow us to predict the turning point $x_{\min}$ for $p_{0}\neq-\pi/2$ ,
@@
Figure. 2 shows how the turning point varies with different $p_{0}$ values at different $\sigma$. The numerical and analytical turning points are extremely well lined up, solidifying the validity of semiclassical calculations. Moreover, the equation suggests that for $x_{\min}$ to be zero, the two hopping $t_{1}$ and $t_{2}$ must be equal, leading to a more simplified expression $x_{\min}=x_{0}(\cos p_{0})^{\frac{1}{\sigma}}$, when $p_{0}=\pm\pi/2$, only then $x_{\min}$ becomes 0. Hence, in the case of $t_{1}\neq t_{2}$, the wave packet has no initial momentum $p_{0}$ value for which the turning point becomes 0. It also validates our previous numerical finding, i.e., only when $t_{1}=t_{2}$, the model can show the synthetic horizon effect, and that too for $p_{0}=-\pi/2$, which also has been observed for the tight binding lattice model [40, 38, 31].